Shift Summary, Evening Dec 29, 1999

Herman B. White (hwhite@fnal.gov)
Thu, 30 Dec 1999 00:04:55 -0600 (CST)

Shift Summary for Eve shift, 29/Dec/99

On Shift: J. Hamm, H. White

Hours of beam available at all intensities: 8.25
Hours of beam available at nominal intensity: 8.00
Hours of E799 data (special runs not included): 7.77
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
Ktevana errors, pipeline reset, tape change and run starts.

========================================================================
Runs ended this shift:

Run 15317

Tapes brought to FCC this shift:
(including tapes for online split)

UP(A,B,C) 283 - 285
UPS095
UPLA13
UPX018
UPF069
UPE019
UPH051

========================================================================
Accesses during shift:

No access

========================================================================
Overview of shift:

Inherited experiment during Run# 15317. Run ended with full
tapes allowing Hogan to raise DC1 voltages by 30 volts, and
Jason to remove 0.5ns of RF delay to ET system to reduce drift in beam
with respect to RF timing.
Work was done on spill quality affecting the live time.

========================================================================
Description of Shift:

15317 15:15 new shift

19:30 Onsplit request Ktev4 UPH052
19:50 Tapes full Run 15317 ends

19:51 Between runs, JH removes .5ns of RF delay to ET,
countering drift in beam wrt Rf timing.

20:00 Hogan readjust DC1HV (window) higher by 30 volts
effective after run 15317. This action is expected to
improve the overall track-link efficiency.

15318 20:04 Start run 15318
ET is not loaded properly for the 1st three spills.
20:05 Online split requests
UPLA13 on Ktev3
UPF070 on Ktev2
UPM052 on Ktev1
20:55 Noticed that livetime dropped to ~55% with large 2Hz
oscillations. MCR informed and elect to reduce intensity
for about 5 mins to investigate the problem.
21:15 MCR changed from 3 turns to 2 turns earlier to reduce
radiation losses. After returning to 3 turns, live time and
stability improves. Cause is still unknown.
22:25 We receive a welcomed visit from the MCR crew chief(Salah)
He wishes to see first hand the monitoring displays. We
print a copy of the spill livetime plot, explain same, and
answered questions. He will include this plot in the MCR
log book as reference for the next shift. They are unsure
what changes caused our livetime drop.
23:30 Channels 1508 and 1512 drop out of laser illumination at
spill 152; fixed by a pipeline reset.


========================================================================
Follow up Notes on Previous Shift
Dave N. still searching for electrical drawings for the gas house.

========================================================================
Items to Follow up on Next Shift

Still monitoring rack fan in RR161 second nim crate from top. There is a
noticable virbration to the touch.

--------------------------------------------------------